14) The rate constant of a first-order reaction is 3.00 × 10^−4 s^−1 at 350.°C. If the activation energy is 149 kJ/mol, calculate the temperature at which its rate constant is 7.60 × 10^−4 s^−1.

_____ *C
Answer #1

Given

k1 = 3.00 \times 10-4 s-1

k2 = 7.60 \times 10-4 s-1

T1 = 350 0C = 350 + 273.15 = 623.15 K

T2 = ?

Ea= 149 kJ/mol = 149000 J/mol

R = 8.314 J/mol.K

Now we have the formula

ln(k2 /k1) = -(Ea/R).(1/T2 - 1/T1 )

by rearranging we get

-Rln(k2 /k1)/ Ea = 1/T2 - 1/T1

-Rln(k2 /k1)/ Ea + 1/T1 = 1/T2

T2 = 1 / {1/T1 - Rln(k2 /k1)/ Ea}

T2 = EaT1 / {Ea - T1R.ln(k2 /k1) }

so by using above values and formula we get

T2 = 149000 J/mol \times 623.15 K / { 149000 J/mol - 623.15 K \times 8.314 J/mol.K \times (ln 7.60 \times 10-4 s-1/3.00 \times 10-4 s-1 )

= 92849350 J.K/mol / { 149000 J/mol - 3590.86)

= 638.53 K

= 638.53 - 273.15

= 365.38 0C

So our asnwer is 365 0C